# SADA: Safe and Adaptive Inference with Multiple Black-Box Predictions

### v1

@@ 5.2 Benchmark datasets
6 Conclusions
To wrap up, we propose a novel algorithm that safely and adaptively aggregates multiple predictions generated by different state-of-the-art ML models. Our method guarantees improved efficiency over the naive approach and demonstrates data adaptivity in balancing predictions of varying quality. Our method can be extended to the situations under distribution shift. In those settings, developing methods that are robust to distribution shift is essential for enhancing the reliability and practical effectiveness of semi-supervised learning.
Appendix A Semiparametric Efficiency Bound
@@
where the zero-mean function $\bm{\phi}(\cdot)$ is called the influence function of $\widehat{\bm{\beta}}$. Then, the central limit theorem implies that $\sqrt{n}(\widehat{\bm{\beta}}-\bm{\beta})\xrightarrow{d}\mathcal{N}(\mathbf{0},\mathbb{E}(\bm{\phi}\bm{\phi}^{\mathrm{T}}))$, provided that $\mathbb{E}(\bm{\phi}\bm{\phi}^{\mathrm{T}})$ is finite and nonsingular. Among all RAL estimators for $\bm{\beta}$, the influence function of the one with the smallest asymptotic variance is called the efficient influence function (EIF), $\bm{\phi}_{\text{eif}}$, and the semiparametric efficiency bound is $\mathbb{E}\left(\bm{\bm{\phi}}_{\text{eif}}\bm{\bm{\phi}}_{\text{eif}}^{\mathrm{T}}\right)$.
@@
We refer readers to BickelKlaassenRitovWellner1993 and Tsiatis2006 for further interpretations. In addition, Kennedy2023 provided several strategies for deriving the efficient influence function together with many examples.
Appendix B Procedure of Inference
The asymptotic normality of $\widehat{\bm{\theta}}^{\text{sada}}$ established in Theorem 1 enables us to construct confidence intervals for $\bm{\theta}$$^{*}$ following standard procedure. Specifically, let
$\widehat{\mathbf{H}}=n^{-1}\sum_{i=1}^{n}\partial\mathbf{s}(\mathbf{X}_{i},Y_{i};\widehat{\bm{\theta}}^{\text{sada}})/\partial{\bm{\theta}}$, and $\widehat{\bm{\Sigma}}_{\text{opt}}=\widehat{\bm{\Sigma}}_{\text{nv}}-(N-n)/N\cdot\widehat{\bm{\Sigma}}_{g}$, where
@@ Appendix B Procedure of Inference
where $\mathcal{R}_{\alpha}$ is a region in $\mathbb{R}^{p}$ such that $\mathbb{P}(\bm{Z}_{p}\in\mathcal{R}_{\alpha})=1-\alpha$ for a $p$-dimensional standard normal random vector $\mathbf{Z}_{p}$. For inference of $j$th component of $\bm{\theta}$$^{*}$, $\theta_{j}^{*}$, let $\widehat{\bm{\Omega}}_{jj}$ be the $j$th diagonal entry of $\widehat{\bm{\Omega}}$. Then, a confidence interval $C_{j,\alpha}$ of $\theta_{j}^{*}$ with error level $\alpha$ is given by
@@
Appendix C Proofs for the Mean Estimation
We provide the proofs for the two cases of the adaptivity properties described in Section 3.
Proof of case 1 (the oracle case).
By the inversion formula of $2\times 2$ block matrix (LuShiou2002, Theorem 2.1), we have
$\operatorname{var}(\widehat{\mathbf{Y}})^{-1}=\left(\begin{matrix}C_{11}^{-1}+C_{11}^{-1}\bm{C}_{21}\bm{M}^{-1}\bm{C}_{21}^{\mathrm{\scriptscriptstyle T}}C_{11}^{-1}&-C_{11}^{-1}\bm{C}_{21}\bm{M}^{-1}\\ -\bm{M}^{-1}\bm{C}_{21}^{\mathrm{\scriptscriptstyle T}}C_{11}^{-1}&\bm{M}^{-1}\end{matrix}\right),$ (C.1)
$\mathbb{E}\{(\widehat{\theta}^{\text{sada}}-\theta^{*})^{2}]=\frac{1}{N}\operatorname{var}(Y).$
Proof of case 2 (Proposition 1).
We derive the EIF for the general estimating equation (6) in the proof of Theorem 2(ii). Since the mean estimation is a special case of (6), we omit its proof here.
and
Similar to (C.1), by the inversion formula of $2\times 2$ block matrix, we obtain
where $n/N\to\pi$ and
Appendix D Proof of Proposition 2
Recall $\mathcal{W}=(\mathcal{W}_{1}^{\mathrm{\scriptscriptstyle T}},\mathcal{W}_{2}^{\mathrm{\scriptscriptstyle T}},\dots,\mathcal{W}_{K}^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}\in\mathbb{R}^{(Kp)\times p}$, $\mathcal{S}(\mathbf{x},\widehat{\mathbf{y}};\theta)=(\mathbf{s}(\mathbf{x},\widehat{y}_{1};\theta)^{\mathrm{\scriptscriptstyle T}},\cdots,\mathbf{s}(\mathbf{x},\widehat{y}_{K};\theta)^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$, and $\widehat{\bm{\theta}}(\mathcal{W})$ solves
$\frac{1}{n}\sum_{i=1}^{n}\mathbf{s}(\mathbf{x}_{i},y_{i};{\bm{\theta}})+\mathcal{W}^{\mathrm{\scriptscriptstyle T}}\left\{\frac{1}{N-n}\sum_{i=n+1}^{N}\mathcal{S}(\mathbf{x}_{i},\widehat{\mathbf{y}}_{i};{\bm{\theta}})-\frac{1}{n}\sum_{i=1}^{n}\mathcal{S}(\mathbf{x}_{i},\widehat{\mathbf{y}}_{i};{\bm{\theta}})\right\}=\mathbf{0}.$
Then the standard Taylor expansion yields that
@@ Appendix D Proof of Proposition 2
where $\overline{{\bm{\theta}}}$ lies between $\widehat{\bm{\theta}}(\mathcal{W})$ and $\bm{\theta}$$^{*}$. Then following uniform weak law of large number (NeweyMcFadden1994) under the regularity conditions, we have
@@
$\mathbb{E}[\{\widehat{\bm{\theta}}(\mathcal{W})-{\bm{\theta}}^{*}\}^{\otimes 2}]\doteq$ $~\mathbf{H}^{-1}\Bigg[\frac{1}{n}\operatorname{var}\{\mathbf{s}(\mathbf{X},Y;{\bm{\theta}}^{*})\}+\frac{N}{n(N-n)}\mathcal{W}^{\mathrm{\scriptscriptstyle T}}\operatorname{var}\{\mathcal{S}(\mathbf{X},\widehat{\mathbf{Y}};{\bm{\theta}}^{*})\}\mathcal{W}$
@@
$\mathbb{E}[\{\widehat{\bm{\theta}}(\mathcal{W}^{\text{opt}})-{\bm{\theta}}^{*}\}^{\otimes 2}]\preceq\mathbb{E}[\{\widehat{\bm{\theta}}(\mathcal{W})-{\bm{\theta}}^{*}\}^{\otimes 2}],~~\forall\mathcal{W}\in\mathbb{R}^{(Kp)\times p},$ (D.1)
@@
Therefore, (D.1) holds by definition.
Appendix E Proof of Theorem 1
Following arguments similar to those in the proof of Proposition 2 in Appendix D, under assumption 1, we have
$\widehat{\bm{\theta}}^{\text{sada}}-{\bm{\theta}}^{*}$
@@ Appendix E Proof of Theorem 1
$\quad-\mathbf{H}^{-1}(\widehat{\mathcal{W}}^{\text{opt}}-\mathcal{W}^{\text{opt}})^{\mathrm{\scriptscriptstyle T}}\left\{\frac{1}{N-n}\sum_{i=n+1}^{N}\mathcal{S}(\mathbf{x}_{i},\widehat{\mathbf{y}}_{i};{\bm{\theta}}^{*})-\frac{1}{n}\sum_{i=1}^{n}\mathcal{S}(\mathbf{x}_{i},\widehat{\mathbf{y}}_{i};{\bm{\theta}}^{*})\right\}$
$~=:T_{1}+T_{2},$ (E.1)
@@
Appendix F Proof of Theorem 2
Proof of (i).
Without loss of generality, we assume $\widehat{Y}_{1}\equiv Y$. We denote $\mathbf{s}\equiv\mathbf{s}_{1}(\mathbf{X},\widehat{Y}_{1};\theta^{*})\equiv\mathbf{s}(\mathbf{X},Y;\theta^{*})$ and $\mathbf{s}_{-1}\equiv(\mathbf{s}(\mathbf{X},\widehat{Y}_{2};\theta^{*})^{\mathrm{\scriptscriptstyle T}},\cdots,\mathbf{s}(\mathbf{X},\widehat{Y}_{K};\theta^{*})^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$. Then
@@ Appendix F Proof of Theorem 2
and $\mathbb{E}\{\mathcal{S}(\mathbf{X},\widehat{\mathbf{Y}};{\bm{\theta}}^{*})\mathbf{s}(\mathbf{X},Y;{\bm{\theta}}^{*})^{\mathrm{\scriptscriptstyle T}}\}=(\operatorname{var}(\mathbf{s}_{1}),\operatorname{cov}(\mathbf{s}_{-1},\mathbf{s}_{1})^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$. Similar to the proof in Appendix C, by the inversion formula of block matrix, we have $\mathcal{W}^{\text{opt}}=(\mathbf{I},\mathbf{0},\dots,\mathbf{0})^{\mathrm{\scriptscriptstyle T}}\cdot(N-n)/N$, and
@@
Proof of (ii).
We first derive the EIF for estimating $\bm{\theta}$$^{*}$, based on labeled data $\mathcal{L}=\{(\mathbf{x}_{i},y_{i}),i=1,\dots,n\}$ and unlabeled data $\mathcal{U}=\{\mathbf{x}_{i},i=n+1,\dots,N\}$. The joint density from one observation is
@@
Then, by the orthogonal conditions (Tsiatis2006, Theorems 4.2 and 4.3), we can obtain
@@
where $\bm{\mu}(\mathbf{x})=\mathbb{E}\{\mathbf{s}(\mathbf{x},Y;{\bm{\theta}}^{*})\mid\mathbf{x}\}$, $\frac{r}{\pi}\{\mathbf{s}(\mathbf{x},y;{\bm{\theta}}^{*})-\bm{\mu}(\mathbf{x})\}\in\Lambda_{1}$ and $\bm{\mu}(\mathbf{x})\in\Lambda_{2}$.
@@
and $\mathbb{E}\{\mathcal{S}(\mathbf{X},\widehat{\mathbf{Y}};{\bm{\theta}}^{*})\mathbf{s}(\mathbf{X},Y;{\bm{\theta}}^{*})^{\mathrm{\scriptscriptstyle T}}\}=(\operatorname{var}\{\bm{\mu}(\mathbf{X})\},\operatorname{cov}\{\mathbf{s}_{-1}(\mathbf{X}),\bm{\mu}(\mathbf{X})\}^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$. Similar to the proof in Appendix C, by the inversion formula of block matrix,, we have $\mathcal{W}^{\text{opt}}=(\mathbf{I},\mathbf{0},\dots,\mathbf{0})^{\mathrm{\scriptscriptstyle T}}\cdot(N-n)/N$. Then, by (E), we have
$\sqrt{N}\{\widehat{\bm{\theta}}^{\text{sada}}-{\bm{\theta}}^{*}\}=$ $~-\mathbf{H}^{-1}\left[\frac{1}{n}\sum_{i=1}^{n}\mathbf{s}(\mathbf{x}_{i},y_{i};{\bm{\theta}}^{*})+\frac{N-n}{N}\left\{\frac{1}{N-n}\sum_{i=n+1}^{N}\mathbf{s}(\mathbf{x}_{i},\widehat{y}_{1,i};{\bm{\theta}}^{*})-\frac{1}{n}\sum_{i=1}^{n}\mathbf{s}(\mathbf{x}_{i},\widehat{y}_{1,i};{\bm{\theta}}^{*})\right\}\right]$
$=$ $~-\mathbf{H}^{-1}\left[\frac{\sqrt{N}}{n}\sum_{i=1}^{n}\{\mathbf{s}(\mathbf{x}_{i},y_{i};{\bm{\theta}}^{*})-\bm{\mu}(\mathbf{x}_{i})\}+\frac{1}{\sqrt{N}}\sum_{i=1}^{N}\widehat{\mu}(\mathbf{x}_{i})\right]+o_{p}(1)$
$=$ $~\frac{1}{\sqrt{N}}\sum_{i=1}^{N}\bm{\Phi}_{\text{eif}}(r_{i},\mathbf{x}_{i},y_{i})+o_{p}(1),$

### v2

@@ 6.1 Politeness of online requests
6.2 Image Classification
In this section, for prediction purposes, we apply SADA to the ImageNet-5 data set, which is constructed from a random subset of five classes from ImageNet-1k (Russakovsky et al., 2015). The data consist of five visually diverse object categories, including animals, artifacts, and natural scenes, each paired with a ground truth class label from the human annotated ImageNet hierarchy. The data set contains 1300 images per class, for a total of 6500 images across five classes, and is split into 5000 training images, 750 validation images, and 750 test images. Among the training images, we randomly select 2500 as unlabeled data and vary the labeled-unlabeled ratio from 0.2 to 1.
We generate two pseudo labels $\widehat{Y}_{1}$ and $\widehat{Y}_{2}$ by training two teacher models, EfficientNet-B0 (Tan & Le, 2020) and RegNetY-008 (Radosavovic et al., 2020), using the following hyperparameters: 10 training epochs, a learning rate of 1e-3, a batch size of 256, and the AdamW optimizer. These trained models are then used to generate pseudo labels for all 5000 training images.
We evaluate the different methods by training DaViT-T (Ding et al., 2022), a popular vision transformer architecture that achieves state-of-the-art performance on ImageNet. The model is trained for 15 epochs with a learning rate of 1e-3, a batch size of 256, and the default AdamW optimizer. Figure 4 reports the Top-1 accuracy on the test set for the naive baseline, PPI using $\widehat{Y}_{1}$ and $\widehat{Y}_{2}$ separately, and SADA across labeled-unlabeled ratios ranging from 0.2 to 1. It is shown that SADA consistently achieves the best performance in terms of prediction accuracy across all settings.
Figure 4: Top-1 accuracy of various methods on the test images across increasing numbers of labeled data.
7 Discussion
In this paper, we propose a novel procedure, SADA, that safely and adaptively aggregates multiple predicted labels originating from diverse state-of-the-art machine learning models for both inference and prediction purposes. An attractive feature of SADA is that it does not require any assumptions on the form, quality, or source of the predictions, which may range from highly accurate and close to the ground truth to extremely poor and essentially random noise. The algorithm automatically aggregates these predictions by assigning them appropriate weights, leading to more efficient inference or improved predictive accuracy. In addition, SADA is practitioner friendly, as it does not require tuning of hyperparameters or any preliminary assessment of the quality of the available predictions.
There are several directions for future exploration. An essential assumption in this paper is that the labeled and unlabeled data are drawn from the same distribution. In settings with distribution shift, the current method should be adapted accordingly, because the augmentation terms in (3) will not converge to zero. In such cases, distributional adjustment based on the type of shift, such as covariate shift or label shift between the labeled and unlabeled data, is necessary to restore the effectiveness of the proposed methods. We are going to investigate these extensions in future work.
Supplementary material
The supplementary material contains all technical proofs as well as additional experimental details. The R package, sada, for implementing the proposed algorithms is available at https://github.com/jw-shan/sada.
SUPPLEMENT
Appendix S1 Semiparametric efficiency bound
@@ 7 Discussion
where the zero-mean function $\bm{\phi}(\cdot)$ is called the influence function of $\widehat{\bm{\beta}}$. Then, the central limit theorem implies that $\sqrt{n}(\widehat{\bm{\beta}}-\bm{\beta})\xrightarrow{d}\mathcal{N}(\mathbf{0},\mathbb{E}(\bm{\phi}\bm{\phi}^{\mathrm{\scriptscriptstyle T}}))$, provided that $\mathbb{E}(\bm{\phi}\bm{\phi}^{\mathrm{\scriptscriptstyle T}})$ is finite and nonsingular. Among all RAL estimators for $\bm{\beta}$, the influence function of the one with the smallest asymptotic variance is called the efficient influence function (EIF), $\bm{\phi}_{\text{eif}}$, and the semiparametric efficiency bound is $\mathbb{E}\left(\bm{\phi}_{\text{eif}}\bm{\phi}_{\text{eif}}^{\mathrm{\scriptscriptstyle T}}\right)$.
@@
We refer readers to Bickel et al. (1993) and Tsiatis (2006) for further interpretations. In addition, Kennedy (2023) provided several strategies for deriving the efficient influence function together with many examples.
Appendix S2 Procedure of inference
The asymptotic normality of $\widehat{\bm{\theta}}^{\text{sada}}$ established in Theorem 1 enables us to construct confidence intervals for $\bm{\theta}$$^{*}$ following standard procedure. Specifically, let $\widehat{\mathbf{H}}=n^{-1}\sum_{i=1}^{n}\partial\mathbf{s}(\mathbf{X}_{i},Y_{i};\widehat{\bm{\theta}}^{\text{sada}})/\partial{\bm{\theta}}$, $\widehat{\bm{\Sigma}}_{\text{opt}}=\widehat{\bm{\Sigma}}_{\text{nv}}-(N-n)/N\cdot\widehat{\bm{\Sigma}}_{g}$, where
@@
where $\mathcal{R}_{\alpha}$ is a region in $\mathbb{R}^{p}$ such that $\mathbb{P}(\bm{Z}_{p}\in\mathcal{R}_{\alpha})=1-\alpha$ for a $p$-dimensional standard normal random vector $\mathbf{Z}_{p}$. For point-wise inference of $j$th component of $\bm{\theta}$$^{*}$, $\theta_{j}^{*}$, let $\widehat{\bm{\Omega}}_{jj}$ be the $j$th diagonal entry of $\widehat{\bm{\Omega}}$. Then, a confidence interval $C_{j,\alpha}$ of $\bm{\theta}$$^{*}$ with error level $\alpha$ is given by
@@
Appendix S3 Proof of Proposition 1
Proof.
Recall that $\widehat{\bm{\theta}}(\mathcal{W})$ solves
$\frac{1}{n}\sum_{i=1}^{n}\mathbf{s}(\mathbf{x}_{i},y_{i};{\bm{\theta}})+\mathcal{W}^{\mathrm{\scriptscriptstyle T}}\left\{\frac{1}{N-n}\sum_{i=n+1}^{N}\mathcal{S}(\mathbf{x}_{i},\widehat{\mathbf{y}}_{i};{\bm{\theta}})-\frac{1}{n}\sum_{i=1}^{n}\mathcal{S}(\mathbf{x}_{i},\widehat{\mathbf{y}}_{i};{\bm{\theta}})\right\}=\mathbf{0},$
where $\mathcal{W}=(\mathcal{W}_{1}^{\mathrm{\scriptscriptstyle T}},\mathcal{W}_{2}^{\mathrm{\scriptscriptstyle T}},\dots,\mathcal{W}_{K}^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}\in\mathbb{R}^{(Kp)\times p}$, and $\mathcal{S}(\mathbf{x},\widehat{\mathbf{y}};{\bm{\theta}})=(\mathbf{s}(\mathbf{x},\widehat{y}_{1};{\bm{\theta}})^{\mathrm{\scriptscriptstyle T}},\cdots,\mathbf{s}(\mathbf{x},\widehat{y}_{K};{\bm{\theta}})^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$. Then the standard Taylor expansion yields that
@@
where $\overline{{\bm{\theta}}}$ lies between $\widehat{\bm{\theta}}(\mathcal{W})$ and $\bm{\theta}$$^{*}$. Then following uniform weak law of large number (Newey & McFadden 1994), under the regularity conditions, we have
@@
$\mathbb{E}[\{\widehat{\bm{\theta}}(\mathcal{W})$- ${\bm{\theta}}^{*}\}^{\otimes 2}]\doteq\mathbf{H}^{-1}\Bigg[\frac{1}{n}\operatorname{var}\{\mathbf{s}(\mathbf{X},Y;{\bm{\theta}}^{*})\}+\frac{N}{n(N-n)}\mathcal{W}^{\mathrm{\scriptscriptstyle T}}\operatorname{var}\{\mathcal{S}(\mathbf{X},\widehat{\mathbf{Y}};{\bm{\theta}}^{*})\}\mathcal{W}$
@@
$\mathbb{E}[\{\widehat{\bm{\theta}}(\mathcal{W}^{\text{opt}})-{\bm{\theta}}^{*}\}^{\otimes 2}]\preceq\mathbb{E}[\{\widehat{\bm{\theta}}(\mathcal{W})-{\bm{\theta}}^{*}\}^{\otimes 2}],~~\forall\mathcal{W}\in\mathbb{R}^{(Kp)\times p},$ (S1)
@@
Therefore, (S1) holds by definition. ∎
Appendix S4 Proofs of adaptivity for mean estimation
We provide the proofs for the two cases of the adaptivity properties in the illustrative mean estimation example described in Section 3.1.
Proof of the oracle case.
By the inversion formula of $2\times 2$ block matrix (Lu & Shiou 2002, Theorem 2.1), we have
$\operatorname{var}(\widehat{\mathbf{Y}})^{-1}=\left(\begin{matrix}C_{11}^{-1}+C_{11}^{-1}\bm{C}_{21}\bm{M}^{-1}\bm{C}_{21}^{\mathrm{\scriptscriptstyle T}}C_{11}^{-1}&-C_{11}^{-1}\bm{C}_{21}\bm{M}^{-1}\\ -\bm{M}^{-1}\bm{C}_{21}^{\mathrm{\scriptscriptstyle T}}C_{11}^{-1}&\bm{M}^{-1}\end{matrix}\right),$ (S2)
$\mathbb{E}\{(\widehat{\theta}^{\text{sada}}-\theta^{*})^{2}\}=\frac{1}{N}\operatorname{var}(Y).$
Proof of Proposition 2..
The EIF for the outcome mean is a special case of Theorem 2(ii), whose proof is given in Supplementary Section S5, so we omit it here.
and
Similar to (S2), by the inversion formula of $2\times 2$ block matrix, we obtain
where $n/N\to\pi$ and
Appendix S5 Proof of Theorem 1
Proof.
Following arguments similar to those in the proof of Proposition 1 in Appendix S3, under assumption 1, we have
$\widehat{\bm{\theta}}^{\text{sada}}-{\bm{\theta}}^{*}$ (S3)
@@
$\qquad-\mathbf{H}^{-1}(\widehat{\mathcal{W}}^{\text{opt}}-\mathcal{W}^{\text{opt}})^{\mathrm{\scriptscriptstyle T}}\left\{\frac{1}{N-n}\sum_{i=n+1}^{N}\mathcal{S}(\mathbf{x}_{i},\widehat{\mathbf{y}}_{i};{\bm{\theta}}^{*})-\frac{1}{n}\sum_{i=1}^{n}\mathcal{S}(\mathbf{x}_{i},\widehat{\mathbf{y}}_{i};{\bm{\theta}}^{*})\right\}$
$~=:T_{1}+T_{2},$
@@
Appendix S6 Proof of Theorem 2
Proof.
Part (i). Without loss of generality, we assume $\widehat{Y}_{1}\equiv Y$. We denote $\mathbf{s}\equiv\mathbf{s}_{1}(\mathbf{X},\widehat{Y}_{1};{\bm{\theta}}^{*})\equiv\mathbf{s}(\mathbf{X},Y;{\bm{\theta}}^{*})$ and $\mathbf{s}_{-1}\equiv(\mathbf{s}(\mathbf{X},\widehat{Y}_{2};{\bm{\theta}}^{*})^{\mathrm{\scriptscriptstyle T}},\cdots,\mathbf{s}(\mathbf{X},\widehat{Y}_{K};{\bm{\theta}}^{*})^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$. Then
@@
and $\mathbb{E}\{\mathcal{S}(\mathbf{X},\widehat{\mathbf{Y}};{\bm{\theta}}^{*})\mathbf{s}(\mathbf{X},Y;{\bm{\theta}}^{*})^{\mathrm{\scriptscriptstyle T}}\}=(\operatorname{var}(\mathbf{s}_{1}),\operatorname{cov}(\mathbf{s}_{-1},\mathbf{s}_{1})^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$. Similar to the proof in Appendix S4, by the inversion formula of block matrix, we have $\mathcal{W}^{\text{opt}}=(\mathbf{I},\mathbf{0},\dots,\mathbf{0})^{\mathrm{\scriptscriptstyle T}}\cdot(N-n)/N$, and
@@
Part (ii). We first derive the EIF for estimating $\bm{\theta}$$^{*}$, based on labeled data $\mathcal{L}=\{(\mathbf{x}_{i},y_{i}),i=1,\dots,n\}$ and unlabeled data $\mathcal{U}=\{\mathbf{x}_{i},i=n+1,\dots,N\}$. The joint density from one observation is
@@
Let
@@
where $\bm{\mu}(\mathbf{x})=\mathbb{E}\{\mathbf{s}(\mathbf{x},Y;{\bm{\theta}}^{*})\mid\mathbf{x}\}$, $\frac{r}{\pi}\{\mathbf{s}(\mathbf{x},y;{\bm{\theta}}^{*})-\bm{\mu}(\mathbf{x})\}\in\Lambda_{1}$ and $\bm{\mu}(\mathbf{x})\in\Lambda_{2}$. It is easy to check $\bm{\Phi}_{\text{eif}}$ satisfies the orthogonal conditions (Tsiatis 2006, Theorems 4.2 and 4.3) and is thus the efficient influence function.
@@
and $\mathbb{E}\{\mathcal{S}(\mathbf{X},\widehat{\mathbf{Y}};{\bm{\theta}}^{*})\mathbf{s}(\mathbf{X},Y;{\bm{\theta}}^{*})^{\mathrm{\scriptscriptstyle T}}\}=(\operatorname{var}\{\bm{\mu}(\mathbf{X})\},\operatorname{cov}\{\mathbf{s}_{-1}(\mathbf{X}),\bm{\mu}(\mathbf{X})\}^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$. Similar to the proof in Appendix S4, by the inversion formula of block matrix,, we have $\mathcal{W}^{\text{opt}}=(\mathbf{I},\mathbf{0},\dots,\mathbf{0})^{\mathrm{\scriptscriptstyle T}}\cdot(N-n)/N$. Then, by (S3), we have
$\sqrt{N}$ $\{\widehat{\bm{\theta}}^{\text{sada}}-{\bm{\theta}}^{*}\}$
$=$ $-\mathbf{H}^{-1}\left[\frac{1}{n}\sum_{i=1}^{n}\mathbf{s}(\mathbf{x}_{i},y_{i};{\bm{\theta}}^{*})+\frac{N-n}{N}\left\{\frac{1}{N-n}\sum_{i=n+1}^{N}\mathbf{s}(\mathbf{x}_{i},\widehat{y}_{1,i};{\bm{\theta}}^{*})-\frac{1}{n}\sum_{i=1}^{n}\mathbf{s}(\mathbf{x}_{i},\widehat{y}_{1,i};{\bm{\theta}}^{*})\right\}\right]$
$=$ $-\mathbf{H}^{-1}\left[\frac{\sqrt{N}}{n}\sum_{i=1}^{n}\{\mathbf{s}(\mathbf{x}_{i},y_{i};{\bm{\theta}}^{*})-\bm{\mu}(\mathbf{x}_{i})\}+\frac{1}{\sqrt{N}}\sum_{i=1}^{N}\widehat{\mu}(\mathbf{x}_{i})\right]+o_{p}(1)$
$=$ $\frac{1}{\sqrt{N}}\sum_{i=1}^{N}\bm{\Phi}_{\text{eif}}(r_{i},\mathbf{x}_{i},y_{i})+o_{p}(1),$